OpenStudy (anonymous):

solve the equation by completing the square. Round to the nearest hundredth if necessary. 4. x^2 + 9x – 14= 0 (1 point) 10.35, 1.35 10.35, –1.35 –10.35, –1.35 –10.35, 1.35

OpenStudy (anonymous):

same idea.. use the quadratic formula & solve!!

OpenStudy (anonymous):

its the last one!!

OpenStudy (anonymous):

a=1; b=9; c=-14

OpenStudy (anonymous):

\[-9\pm \sqrt{137}/2\]

OpenStudy (anonymous):

To complete the square take one half of the b value. Which gives you 9/2 Now square it to get 81/4 Which gives us \[x ^{2}+9x+81/4-81/4+14=0\] Rearrange \[(x+9/2)^{2}=14+81/4\] Which now gives us \[(x+9/2)^{2}=56/4+81/4\] Which becomes \[(x+9/2)^{2}=137/4\] \[x+9/2=\pm \sqrt{137/4}\]\[x=-9/2\pm 1/2\sqrt{137}\]